OpenStudy (anonymous):

Multiply. 2 5/6 * 4 1/2

OpenStudy (anonymous):

I got first 153/12 then i simplify an got 51/4

OpenStudy (anonymous):

This is the question right? \[2\frac{ 5 }{ 6 } \times 4\frac{1}{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First I would convert the mixed fractions into improper fractions: \[2\frac{5}{6} is the same as \frac{12 + 5}{6} is the same as \frac{17}{6}\] \[4\frac{1}{2} is the same as \frac{8 + 1}{2} is the same as \frac{9}{2}\]

OpenStudy (anonymous):

do you see how I did that? You multiply the number outside the fraction by the denominator(bottom) of the fraction and add it to the numerator(top)

OpenStudy (anonymous):

i got 153/12

OpenStudy (anonymous):

yes that's correct!

OpenStudy (anonymous):

51/4 is the correct reduction as well

OpenStudy (anonymous):

the question is asking me to simplify i got 51/4

OpenStudy (anonymous):

thank you! <3

OpenStudy (anonymous):

no problem